The mean serum level measured in 12 patients twenty-four hours after they received a newly recommended antibiotic was 1.2 mg/dl with a standard deviation of 0.4 mg/dl. If the mean serum level in the general population is 1.0 mg/dl, test whether or not the mean serum level in the sample group is significantly different from that of the general population. Use a = 5%. Compute the test value.


1
Expert's answer
2022-05-25T17:01:20-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=1"

"H_1:\mu\not=1"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.05," and the critical value for a two-tailed test is "z_c = 1.96."

The rejection region for this two-tailed test is "R = \{z:|z|>1.96\}."

The z-statistic is computed as follows:



"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{1.2-1}{0.4/\sqrt{12}}\approx1.732"

Since it is observed that "|z|=1.732<1.96=z_c," it is then concluded that the null hypothesis is not rejected.

Using the P-value approach:

The p-value is "p=2P(z>1.732)= 0.083274," and since "p= 0.083274>0.05=\alpha," it is concluded that the null hypothesis is not rejected.

Therefore, there is not enough evidence to claim that the population mean "\mu"

is different than 1, at the "\alpha = 0.05" significance level.